OpenStudy (anonymous):

Titan orbits Saturn in a circular orbit of radius 1.22 x 109 m and a period of 15.9 days. From this data, determine the mass of Saturn.

OpenStudy (mos1635):

 F=Fcentripetal F=Fgravitation Fc=Fg m*V^2/R=G*M*m/R^2 V^2=GM/R (2πR/T)^2=GM/R ..... T=sqr (4π^2*R^3/GM) G and M of saturn needed